# Joint Node Selection and Resource Allocation Optimization for Cooperative Sensing with a Shared Wireless Backhaul

### v1

@@ III-C Target Localization at the FC
Suppose that the FC can obtain the estimated time delays $\{\hat{\tau}_{n}\}$ and effective reflecting coefficients $\{\hat{\alpha}_{n}\}$ (information-domain) at the receivers with negligible quantization error, then the eigenmatrix $\mathbf{U}$ can be acquired with negligible quantization loss, which implicitly contains the information-domain signals $\{\hat{\tau}_{n},\hat{\alpha}_{n}\}$. After receiving the quantized signal samples $\tilde{\mathbf{r}}_{nC}$ (signal-domain) from the $N$ receivers, the FC is able to recover the target location by fusing these signal-domain and information-domain measurements, i.e., $\{\hat{\tau}_{n},\hat{\alpha}_{n}\}$ and $\{\left[\mathbf{r}_{nC}\right]_{j},j\in[1,2K_{n}]\}$.
@@
Then, by performing inverse linear transformation on $\tilde{\mathbf{r}}_{nC}$, i.e., multiplying $\tilde{\mathbf{r}}_{nC}$ on the left by eigenmatrix $\mathbf{U}$, we are able to recover the received signal samples by
$\tilde{\mathbf{r}}_{n}=\mathbf{U}\tilde{\mathbf{r}}_{nC}=\mathbf{r}_{n}+\mathbf{U}\mathbf{q}_{n},$ (17)
where $\mathbf{U}\mathbf{q}_{n}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{n})$ is the quantization error vector of $\mathbf{r}_{n}$ and $\mathbf{Q}_{n}$ is given by
$\mathbf{Q}_{n}=\mathbf{U}\textrm{diag}([\eta_{n1},\cdots,\eta_{n(2K_{n})}])\mathbf{U}^{T}.$ (18)
Finally, the FC estimates the target location by resorting to the ML rule. Let $\Omega\subseteq\left\{1,2,\cdots,N\right\}$ denote the selected receivers (the detailed selection strategy will be introduced in Section IV), then the received signal samples available at the FC can be represented as $\tilde{\mathbf{r}}_{\Omega}=\left\{\tilde{\mathbf{r}}_{n}\mid n\in\Omega\right\}$. Therefore, the ML estimation of $\boldsymbol{\theta}=[x,y]^{T}$ can be obtained by
$\hat{\boldsymbol{\theta}}=\arg\max_{\boldsymbol{\theta}}\ln p(\tilde{\mathbf{r}}_{\Omega}\mid\boldsymbol{\theta}),$ (19)
@@
$\begin{split}&\ln p(\tilde{\mathbf{r}}_{\Omega}\mid\boldsymbol{\theta})=\\ &-\frac{1}{2}\sum_{n\in\Omega}\left[(\tilde{\mathbf{r}}_{n}-\mathbf{s}_{n})^{T}\left(\mathbf{Q}_{\omega_{n}}+\mathbf{Q}_{n}\right)^{-1}(\tilde{\mathbf{r}}_{n}-\mathbf{s}_{n})\right]+D_{2},\end{split}$ (20)
and $D_{2}$ is a constant uncorrelated with $\boldsymbol{\theta}$, $\mathbf{Q}_{\omega_{n}}=\frac{1}{2}\sigma^{2}_{n}\mathbf{I}_{2K_{n}}$ is a diagonal matrix and $\mathbf{s}_{n}=[\{\sqrt[]{E}\Re(\hat{\alpha}_{n})s(k_{n}T_{s}-\tau_{n}),k_{n}\in\mathcal{K}_{n}\},$${E}\Im(\hat{\alpha}_{n})s(k_{n}T_{s}-\tau_{n}),k_{n}\in\mathcal{K}_{n}\}]^{T}$.
@@ IV-A Optimization Problem Formulation
Then, according to the MAC channel capacity limits [33], the number of quantization bits $\{X_{nj}\}$ must satisfy the following conditions for each possible non-empty subset $\mathcal{S}$ of $\Omega$:
$\sum_{n\in\mathcal{S}}\frac{\sum_{j=1}^{2K_{n}}X_{nj}}{W}\leq\log\left(1+\frac{\sum_{n\in\mathcal{S}}P_{n}g_{n}}{N_{0}}\right),$ (23)
@@
where $W$ is the number of channel uses, $N_{0}$ denotes the noise power, $P_{n}$ and $g_{n}$ represent the transmit power at the $n$-th sensing receiver and the channel gain between the $n$-th receiver and the FC, respectively. Therefore, the considered optimization problem can be formulated as [^1] Note that in this work, we assume that the $N$ receivers transmit signals using the same power, thus the transmit power $P_{n}$ is not an optimization variable.
@@ IV-B1 MCSCA Algorithm for Quantization Bit Allocation
Next, since the eigenmatrix $\mathbf{U}$ is an orthogonal matrix, $\mathbf{J}(\hat{\boldsymbol{\theta}})$ can be equivalently rewritten as
@@
$\sum_{n\in\Omega}\frac{\partial\mathbf{s}_{n}}{\partial\hat{\boldsymbol{\theta}}}\mathbf{U}\operatorname{diag}\left([y_{n}(X_{n1}),\cdots,y_{n}(X_{n(2K_{n})})]\right)(\frac{\partial\mathbf{s}_{n}}{\partial\hat{\boldsymbol{\theta}}}\mathbf{U})^{T},$
@@ IV-B2 Convergence Analysis for the MCSCA Algorithm
Please refer to Appendix A for the proof. ∎
@@ V Simulation Results
For comparison, the following three baselines are considered: 1) A TOA based IDCS scheme, where each receiver only sends the estimated time delay $\hat{\tau}_{n}$ and effective reflecting coefficient $\hat{\alpha}_{n}$ to the FC for target localization based on the ML rule; 2) A uniform-quantization SDCS scheme, where each echo signal sample is uniformly quantized using 8 bits and then transmitted to the FC for localization; 3) An ideal SDCS scheme with unlimited communication capacity, i.e., assuming perfect echo signals are available at the FC.
Figure 3: Convergence behavior of the proposed MCSCA algorithm
@@
Figure 4: Comparison between the proposed and baseline designs under different SNR regimes
Then, we compare in Fig. 4 the MSE performance and the average channel use number $W$ between the baseline designs and the proposed HISDCS design under different SNR regimes. It can be observed that in the medium-to-low SNR regime, the proposed HISDCS scheme outperforms the IDCS design in terms of MSE performance. Besides, it is worth noting that employing a smaller $\epsilon$ in constraint (24b) enhances the MSE performance of the HISDCS design. Especially when $\epsilon$ is set to $1.01\times\epsilon^{*}$ ($\epsilon^{*}$ denotes the minimum achievable CRLB as mentioned in Section IV), the proposed HISDCS scheme outperforms the uniform-quantization SDCS scheme and its performance is remarkably close to the ideal SDCS design, which is regarded as the localization performance lower bound. Besides, from Fig. 4 (b), it is evident that the proposed HISDCS scheme sigificantly reduces the communication overhead as compared to the uniform-quantization SDCS scheme, since the former only needs 8 channel uses, while almost 101 channel uses are required by the latter.
@@
Figure 5: Comparison between the proposed and baseline designs under different numbers of sensing receivers
Fig. 5 presents the MSE performance and average channel use number comparison between the baseline designs and the proposed HISDCS design under different numbers of sensing receivers $N$, where we set the average SNR to 0 dB. From Fig. 5 (a), it is seen that the localization performance of the proposed HISDCS scheme is superior to that of the IDCS scheme. Additionally, as $N$ increases, the MSE of all the considered schemes decreases, which is reasonable since larger $N$ implies higher cooperation gain which is beneficial for localization performance improvement. Besides, from Fig. 5 (b), it can be observed that the communication overhead of the proposed HISDCS scheme is substantially reduced as compared to that required by the uniform-quantization SDCS scheme, and with the increasing of $N$, there is a consistent decrease in $W$ across all schemes, which is mainly due to the fact that the amount of information transmitted to the FC for localization performance improvement increases as $N$ increases.
Figure 6: Channel use number comparison under different SNR regimes
@@
Figure 7: MSE performance comparison under different topologies
@@
In Fig. 6, we show the average number of channel uses $W$ required by the proposed HISDCS scheme with or without node selection, and compare the performance of the proposed node selection strategy with the low-complexity bit reallocation algorithm presented in Section IV. From this figure, it is seen that compared to the scheme with the same number of quantization bits allocated to each receiver, the proposed MCSCA algorithm is much better in terms of communication cost. Besides, it can be observed that when node selection is employed, the number of channel uses $W$ can be substantially reduced, for example, $W$ can be reduced by $29.9\%$ when $\textrm{SNR}=0$ dB, indicating a reduction in communication overhead. Moreover, it is noteworthy that the proposed greedy bit reallocation algorithm requires a slightly higher average number of channel use than the scheme with greedy node selection strategy, but it achieves a substantial reduction in computational complexity (as discussed in Section IV).
@@
Figure 8: MSE performance and required channel use number versus the number of sampling points $K_{n}$
Then, we show in Fig. 8 the mean squared error (MSE) performance for target location estimation and the average channel use number $W$ achieved by Algorithm 2 under different numbers of sampling points $K_{n}$. In this simulation, average signal-noise ratio (SNR) of the receivers is set to be -2 dB. From Fig. 4 (a), it can be seen that the localization performance improves with the increasing of $K_{n}$. However, when $K_{n}$ is large, the localization performance improvement gradually saturates, and when $K_{n}$ exceeds 14, the localization performance reaches a plateau. Besides, we can observe from Fig. 8 that the required number of channel uses $W$ first increases as $K_{n}$ increases, and then gradually stabilizes as $K_{n}$ surpasses 4. Hence, the value of $K_{n}$ needs to be carefully designed, for example, setting it to 10, which enables the achievement of favorable localization performance without incurring significant communication overhead, and the complexity of the MCSCA will not be excessively high.
Figure 9: MSE performance and required channel use number versus the sampling frequency $f_{s}$
Fig. 9 illustrates the impact of the sampling frequency $f_{s}$ on the MSE performance and the average channel use number $W$, under the given sampling point number $K_{n}=10$. It is observed that the best localization performance is achieved when $f_{s}$ is about 100 MHz, otherwise, the performance gradually deteriorates. This is mainly due to the fact that, an excessively small or large $f_{s}$ may result in the inability to extract useful localization information within the main lobe region of the echo signal under given $K_{n}$, consequently leading to a decline in the localization performance. Besides, we can see that the channel use number $W$ barely changes as $f_{s}$ varies, which is expected since $K_{n}$ is fixed in this simulation.
Figure 10: MSE performance and the number of channel use comparison under different values of $\epsilon$
Finally, we present in Fig. 10 the MSE performance and the average number of channel use $W$ achieved by the proposed HISDCS scheme under different values of $\epsilon$. It can be observed that as $\epsilon$ increases, the MSE gradually increases while $W$ decreases. This is mainly owing to the fact that the demand for meeting the CRLB constraint gradually diminishes as $\epsilon$ increases, thus leading to localization performance degradation and communication overhead reduction. Hence, there is a tradeoff between these two performance metrics that should be carefully considered based on the practical requirements.
VI Conclusion

### v2

@@ III-C Target Localization at the FC
Suppose that the FC can obtain the estimated time delays $\{\hat{\tau}_{n}\}$ and effective reflecting coefficients $\{\hat{\alpha}_{n}\}$ (information-domain) at the receivers with negligible quantization error, then the eigenmatrices $\{\mathbf{U}_{n}\}$ can be acquired with negligible quantization loss, which implicitly contains the information-domain signals $\{\hat{\tau}_{n},\hat{\alpha}_{n}\}$. After receiving the quantized signal samples $\tilde{\mathbf{r}}_{nC}$ (signal-domain) from the $N$ receivers, the FC is able to recover the target location by fusing these signal-domain and information-domain measurements, i.e., $\{\hat{\tau}_{n},\hat{\alpha}_{n}\}$ and $\{\left[\tilde{\mathbf{r}}_{nC}\right]_{j},j\in[1,2K_{n}]\}$.
@@
Then, by performing inverse linear transformation on $\tilde{\mathbf{r}}_{nC}$, i.e., multiplying $\tilde{\mathbf{r}}_{nC}$ on the left by eigenmatrix $\mathbf{U}_{n}$, we are able to recover the received signal samples by
$\tilde{\mathbf{r}}_{n}=\mathbf{U}_{n}\tilde{\mathbf{r}}_{nC}=\mathbf{r}_{n}+\mathbf{U}_{n}\mathbf{q}_{n},$ (26)
where $\mathbf{U}_{n}\mathbf{q}_{n}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{n})$ is the quantization error vector of $\mathbf{r}_{n}$ and $\mathbf{Q}_{n}$ is given by
$\mathbf{Q}_{n}=\mathbf{U}_{n}\textrm{diag}([\eta_{n1},\cdots,\eta_{n(2K_{n})}])\mathbf{U}_{n}^{T}.$ (27)
Finally, the FC estimates the target location by resorting to the ML rule. Let $\Omega\subseteq\left\{1,2,\cdots,N\right\}$ denote the selected receivers (the detailed selection strategy will be introduced in Section IV), then the received signal samples available at the FC can be represented as $\tilde{\mathbf{r}}_{\Omega}=\left\{\tilde{\mathbf{r}}_{n}\mid n\in\Omega\right\}$. Since the noises $\omega_{n}(t)$ are uncorrelated across the sensing receivers, as noted in Section II. The ML estimation of $\boldsymbol{\theta}\triangleq[x,y]^{T}$ can be obtained by integrating both information-domain and signal-domains as follows:
$\hat{\boldsymbol{\theta}}=\arg\max_{\boldsymbol{\theta}}\ln p(\tilde{\mathbf{r}}_{\Omega}\mid\boldsymbol{\theta},\hat{\boldsymbol{\tau}},\hat{\boldsymbol{\alpha}}),$ (28)
@@
$\begin{split}&\ln p(\tilde{\mathbf{r}}_{\Omega}\mid\boldsymbol{\theta},\hat{\boldsymbol{\tau}},\hat{\boldsymbol{\alpha}})=\\ &-\frac{1}{2}\sum_{n\in\Omega}\left[(\tilde{\mathbf{r}}_{n}-\mathbf{s}_{n})^{T}\left(\mathbf{Q}_{\omega_{n}}+\mathbf{Q}_{n}\right)^{-1}(\tilde{\mathbf{r}}_{n}-\mathbf{s}_{n})\right]+D_{2},\end{split}$ (29)
and $D_{2}$ is a constant uncorrelated with $\boldsymbol{\theta}$, $\mathbf{Q}_{\omega_{n}}\triangleq\frac{1}{2}\sigma^{2}_{n}\mathbf{I}_{2K_{n}}$ is a diagonal matrix, $\hat{\boldsymbol{\tau}}\triangleq\{\hat{\tau}_{n}\mid n=1,\cdots,N\}$, $\hat{\boldsymbol{\alpha}}\triangleq\{\hat{\alpha}_{n}\mid n=1,\cdots,N\}$ and $\mathbf{s}_{n}\triangleq[\{\sqrt[]{E}\Re(\hat{\alpha}_{n}s(k_{n}T_{s}-\tau_{n})),k_{n}\in\mathcal{K}_{n}\},$${E}\Im(\hat{\alpha}_{n}s(k_{n}T_{s}-\tau_{n})),k_{n}\in\mathcal{K}_{n}\}]^{T}$.
Remark 1.
In (28) and (29), it seems that only signal-domain information $\tilde{\mathbf{r}}_{\Omega}$ is used for localization, however, it is important to mention that the information-domain parameters also play a significant role in the overall localization process. Specifically, the term $\mathbf{Q}_{n}$ in (29) is derived from (27), while the eigenmatrix $\mathbf{U}_{n}$ in (27) is obtained by performing an eigenvalue decomposition of the matrix $\mathbf{Q}_{\mathbf{r}_{n}}=\mathbf{U}_{n}\mathbf{\Lambda}_{n}\mathbf{U}_{n}^{T}$, which is computed based on the estimated time delays $\hat{\tau}_{n}$ and reflection coefficients $\hat{\alpha}_{n}$ as described in (21). Besides, the term $\mathbf{s}_{n}$ in (29) also involves $\hat{\alpha}_{n}$ and $\hat{\tau}_{n}$, where $\hat{\alpha}_{n}$ affects the amplitude of $\mathbf{s}_{n}$ and $\hat{\tau}_{n}$ is implicitly entailed in the set $\mathcal{K}_{n}$. Thus, although the final position estimation relies on the quantized signals, the underlying information-domain parameters (time delays and reflection coefficients) are also involved in localization and they are crucial for shaping the quantized signals. Besides, we recognize that although we have integrated information-domain measurements into the signal-domain processing, the proposed method may not be able to perfectly exploit them. Investigating more efficient HISDCS schemes would be an interesting future research direction.
@@ IV-A Optimization Problem Formulation
Then, according to the Gaussian MAC capacity limits [38], the number of quantization bits $\{X_{nj}\}$ must satisfy the following conditions for each possible non-empty subset $\mathcal{S}$ of $\Omega$:
$\sum_{n\in\mathcal{S}}R_{n}=\sum_{n\in\mathcal{S}}\frac{\sum_{j=1}^{2K_{n}}X_{nj}}{W}\leq\log\left(1+\frac{\sum_{n\in\mathcal{S}}P_{n}g_{n}}{N_{0}}\right),$ (32)
@@
where $R_{n}=\frac{\sum_{j=1}^{2K_{n}}X_{nj}}{W}$ represents the transmission rate in the unit of bit per channel use (bpcu), $W$ is the number of channel uses [38], $N_{0}$ denotes the noise power, $P_{n}$ and $g_{n}$ represent the transmit power at the $n$-th sensing receiver and the channel gain between the $n$-th receiver and the FC, respectively. Please refer to Appendix A for the derivation of (32). Therefore, the considered optimization problem can be formulated as [^1] Note that in this work, we assume that the $N$ receivers transmit signals using the same power, thus the transmit power $P_{n}$ is not an optimization variable.
@@ IV-B1 MCSCA Algorithm for Quantization Bit Allocation
Next, since the eigenmatrix $\mathbf{U}_{n}$ is an orthogonal matrix, $\mathbf{J}(\hat{\boldsymbol{\theta}})$ can be equivalently rewritten as
@@
$\sum_{n\in\Omega}\frac{\partial\mathbf{s}_{n}}{\partial\hat{\boldsymbol{\theta}}}\mathbf{U}_{n}\operatorname{diag}\left([y_{n}(X_{n1}),\cdots,y_{n}(X_{n(2K_{n})})]\right)(\frac{\partial\mathbf{s}_{n}}{\partial\hat{\boldsymbol{\theta}}}\mathbf{U}_{n})^{T},$
@@ IV-B2 Convergence Analysis for the MCSCA Algorithm
Please refer to Appendix B for the proof. ∎
@@ IV-B3 Node Selection Strategy
It is important to note that the overall localization performance is influenced not only by the number of quantization bits but also by the network topology. In the proposed algorithm, the geometric configuration of the sensing nodes is taken into consideration through the calculation of the CRLB as described in (30) and (31).
@@
To summarize, although the greedy approach of removing nodes with the fewest allocated bits is not optimal, we believe it offers a reasonable balance between performance and complexity within the current scope of our work.
@@ V Simulation Results
Figure 4: Convergence behavior of the proposed MCSCA algorithm.
@@
Figure 5: Comparison between the proposed and baseline designs under different SNR regimes.
For comparison, the following three baselines are considered: 1) A TOA based IDCS scheme, where each receiver only sends the estimated time delay $\hat{\tau}_{n}$ and effective reflecting coefficient $\hat{\alpha}_{n}$ to the FC for target localization based on the ML rule; 2) A TOA-RSS based IDCS scheme, where the estimated time delay $\hat{\tau}_{n}$ , effective reflecting coefficient $\hat{\alpha}_{n}$ , and the signal power around $\hat{\tau}_{n}$ are transmitted to the FC for localization; 3) A uniform-quantization SDCS scheme, where each echo signal sample is uniformly quantized using 8 bits and then transmitted to the FC for localization; 4) An ideal SDCS scheme with unlimited communication capacity, i.e., assuming perfect echo signals are available at the FC.
@@
Figure 6: Comparison between the proposed and baseline designs under different numbers of sensing receivers.
Then, we compare in Fig. 5 the MSE performance and the average channel use number $\overline{W}$ between the baseline designs and the proposed HISDCS design under different SNR regimes. It can be observed that in the medium-to-low SNR regime, the proposed HISDCS scheme outperforms the TOA based IDCS and the TOA-RSS based IDCS design in terms of MSE performance. This is because incorporating signal-domain information provides a performance boost compared to using only information-domain estimates. Furthermore, even with a limited number of quantization bits for the signal-domain data, the constraint on the CRLB, i.e., $\text{CRLB}_{\hat{\boldsymbol{\theta}}}\leq\epsilon$ , ensures an efficient quantizzation bit allocation scheme that makes the proposed method outperform the IDCS approach. Besides, it is worth noting that employing a smaller $\epsilon$ in constraint (33b) enhances the MSE performance of the HISDCS design. Especially when $\epsilon$ is set to $1.01\times\epsilon^{*}$ ($\epsilon^{*}$ denotes the minimum achievable CRLB as mentioned in Section IV), the proposed HISDCS scheme outperforms the uniform-quantization SDCS scheme and its performance is remarkably close to the ideal SDCS design, which is regarded as the localization performance lower bound. Besides, from Fig. 4 (b), it is evident that the proposed HISDCS scheme sigificantly reduces the communication overhead as compared to the uniform-quantization SDCS scheme, since the former only needs 8 channel uses, while almost 101 channel uses are required by the latter.
@@
Figure 7: MSE performance comparison under a large number of sensing nodes.
Fig. 6 presents the MSE performance and average channel use number comparison between the baseline designs and the proposed HISDCS design under different numbers of sensing receivers $N$, where we set the average SNR to 0 dB. From Fig. 5 (a), it is seen that the localization performance of the proposed HISDCS scheme is superior to that of the TOA based IDCS and TOA-RSS based schemes. Additionally, as $N$ increases, the MSE of all the considered schemes decreases, which is reasonable since larger $N$ implies higher cooperation gain which is beneficial for localization performance improvement. Besides, from Fig. 5 (b), it can be observed that the communication overhead of the proposed HISDCS scheme is substantially reduced as compared to that required by the uniform-quantization SDCS scheme, and with the increasing of $N$, there is a consistent decrease in $\bar{W}$ across all schemes, which is mainly due to the fact that the amount of information transmitted to the FC for localization performance improvement increases as $N$ increases.
Fig. 7 shows the MSE performance comparison under a larger number of nodes. As can be seen, the localization performance of the proposed HISDCS scheme is still superior to that of the TOA based IDCS scheme when $N\geq 10$. This is mainly because the proposed method combines the advantages offered by information-domain parameters and signal-domain information, and the latter allows for a better extraction of the signal-domain features, and therefore enhances the overall MSE performance. However, as the number of nodes increases, the fusion of the delay estimates becomes more accurate, which effectively reduces the impact of individual estimation errors and thereby limits the additional benefits gained from signal-domain processing in the proposed scheme.
Figure 8: Channel use number comparison under different SNR regimes.
Figure 9: MSE performance comparison under different topologies.
@@
In Fig. 8, we show the average number of channel uses $\overline{W}$ required by the proposed HISDCS scheme with or without node selection, and compare the performance of the proposed node selection strategy with the low-complexity bit reallocation algorithm presented in Section IV. From this figure, it is seen that compared to the scheme with the same number of quantization bits allocated to each receiver, the proposed MCSCA algorithm is much better in terms of communication cost. Besides, it can be observed that when node selection is employed, the number of channel uses $\overline{W}$ can be substantially reduced, for example, $\overline{W}$ can be reduced by $30.1\%$ when $\textrm{SNR}=0$ dB, indicating a reduction in communication overhead. Moreover, it is noteworthy that the proposed greedy bit reallocation algorithm requires a slightly higher average number of channel use than the scheme with greedy node selection strategy, but it achieves a substantial reduction in computational complexity (as discussed in Section IV).
@@
Figure 10: MSE performance and required channel use number versus the number of sampling points $K_{n}$.
Then, we show in Fig. 10 the mean squared error (MSE) performance for target location estimation and the average channel use number $\overline{W}$ achieved by Algorithm 2 under different numbers of sampling points $K_{n}$. In this simulation, average signal-noise ratio (SNR) of the receivers is set to be -2 dB. From Fig. 10, it can be seen that the localization performance improves with the increasing of $K_{n}$. However, when $K_{n}$ is large, the localization performance improvement gradually saturates, and when $K_{n}$ exceeds 14, the localization performance reaches a plateau. Besides, we can observe from Fig. 8 that the required number of channel uses $\overline{W}$ first increases as $K_{n}$ increases, and then gradually stabilizes as $K_{n}$ surpasses 4. Hence, the value of $K_{n}$ needs to be carefully designed, for example, setting it to 10, which enables the achievement of favorable localization performance without incurring significant communication overhead, and the complexity of the MCSCA will not be excessively high.
Figure 11: MSE performance and required channel use number versus the sampling frequency $f_{s}$.
Fig. 11 illustrates the impact of the sampling frequency $f_{s}$ on the MSE performance and the average channel use number $\overline{W}$, under the given sampling point number $K_{n}=10$. It is observed that the best localization performance is achieved when $f_{s}$ is about 100 MHz, otherwise, the performance gradually deteriorates. This is mainly due to the fact that, an excessively small or large $f_{s}$ may result in the inability to extract useful localization information within the main lobe region of the echo signal under given $K_{n}$, consequently leading to a decline in the localization performance. Besides, we can see that the channel use number $\overline{W}$ barely changes as $f_{s}$ varies, which is expected since $K_{n}$ is fixed in this simulation.
Figure 12: MSE performance and the number of channel use comparison under different values of $\epsilon$.
Finally, we present in Fig. 12 the MSE performance and the average number of channel use $\overline{W}$ achieved by the proposed HISDCS scheme under different values of $\epsilon$. It can be observed that as $\epsilon$ increases, the MSE gradually increases while $\overline{W}$ decreases. This is mainly owing to the fact that the demand for meeting the CRLB constraint gradually diminishes as $\epsilon$ increases, thus leading to localization performance degradation and communication overhead reduction. Hence, there is a trade-off between these two performance metrics that should be carefully considered based on the practical requirements.
VI Conclusion, Challenges and Future Work